OpenStudy (anonymous):

What is the simplified form of 4x^2+20x+25/2x+5

OpenStudy (saifoo.khan):

\[\frac{(2x+5)(2x+5)}{2x+5}\]

OpenStudy (anonymous):

What is the restriction?

OpenStudy (saifoo.khan):

2x+ 5 answer.

OpenStudy (anonymous):

no canceling? i am disappointed.(ok not really)

OpenStudy (saifoo.khan):

\[\Huge \frac{\cancel{(2x+5)}(2x+5)}{\cancel{2x+5}}\]

OpenStudy (anonymous):

restriction \[x\neq -\frac{5}{2}\]

OpenStudy (anonymous):

saifoo you area a quick study. too damned quick...

OpenStudy (saifoo.khan):

Lolo, i got a i7 processor! =(

OpenStudy (saifoo.khan):

(im my mind)